OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5

OpenStudy (anonymous):

type this into www.myalgebra.com and it'll do it for you.

OpenStudy (anonymous):

okay sure will

OpenStudy (anonymous):

how did you get the answer "No Solution"

OpenStudy (jackellyn):

First you have to distribute -3(x)+(-3)(3)=-3(x)+(-3)(1)-5 Which gives -3x - 9 = -3x -3 -5 Simplify/combine like terms -3x - 9 = -3x -8 Do you know what the next step is?

OpenStudy (anonymous):

No?/:

OpenStudy (jackellyn):

The next step would be to get the x's together. So you add -3x to both sides -3x+3x -9 = -3x+3x -8 Which will cancel the 3x -9 = -8 And that is why there is no solution.

OpenStudy (jackellyn):

Because \[-9 \neq -8\]